OpenStudy (anonymous):

the output of a function is 5 more than 2 times the input. find the input when the output is 17.

OpenStudy (anonymous):

5+2x=17 17-5 = 12 2x=12 12/2 = 6 thus x = 6

OpenStudy (anonymous):

Sorry, I did that wrong, you're right, x=6

OpenStudy (anonymous):

I was so confused as to what you were doing.

OpenStudy (anonymous):

Lol, so was I.

OpenStudy (anonymous):

I was a bit out of it there.